# ECLIPSE: Semantic Entropy-LCS for Cross-Lingual Industrial Log Parsing

### v1

@@ Abstract.
Log parsing, a vital task for interpreting the vast and complex data produced within software architectures faces significant challenges in the transition from academic benchmarks to the industrial domain. Existing log parsers, while highly effective on standardized public datasets, struggle to maintain performance and efficiency when confronted with the sheer scale and diversity of real-world industrial logs. These challenges are two-fold: 1) massive log templates: The performance and efficiency of most existing parsers will be significantly reduced when logs of growing quantities and different lengths; 2) Complex and changeable semantics: Traditional template-matching algorithms cannot accurately match the log templates of complicated industrial logs because they cannot utilize cross-language logs with similar semantics. To address these issues, we propose ECLIPSE, Enhanced Cross-Lingual Industrial log Parsing with Semantic Entropy-LCS, since cross-language logs can robustly parse industrial logs. On the one hand, it integrates two efficient data-driven template-matching algorithms and Faiss indexing. On the other hand, driven by the powerful semantic understanding ability of the Large Language Model (LLM), the semantics of log keywords were accurately extracted, and the retrieval space was effectively reduced. It is worth noting that we launched a Chinese and English cross-platform industrial log parsing benchmark ECLIPSE-Bench to evaluate the performance of mainstream parsers in industrial scenarios. Our experimental results, conducted across public benchmarks and the proprietary ECLIPSE-Bench dataset, underscore the superior performance and robustness of our proposed ECLIPSE. Notably, ECLIPSE both delivers state-of-the-art performance when compared to strong baselines on diverse datasets and preserves a significant edge in processing efficiency [^1] We will release our code and dataset..
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ;
@@ 1. Introduction
Logs play a crucial role in Algorithmic IT Operations (AIOps) behavior compared to other operational and maintenance data, as they provide significant insights into system behavior (He et al., 2021; Zhu et al., 2019). By analyzing logs, we can complete a variety of downstream tasks, such as anomaly detection (Du et al., 2017; Nandi et al., 2016; Zhang et al., 2024a), fault diagnosis (He et al., 2018; Zhang et al., 2024b) and root cause analysis (Tak et al., 2016; Chuah et al., 2010). In general, most log analysis methods make log parsing the primary step in automated log analysis (He et al., 2016). Illustrated in 1, semi-structured logs are generated from log statement code and need to be parsed to structured logs by algorithms. The most challenging task among them is to extract templates from logs (Zhu et al., 2019), which represent the unchanged parts of the log parsing process. Traditional log parsing was accomplished by manually configuring regular expressions (Lang, 2013). With the exponential growth in the volume and complexity of logs due to distributed systems, cloud computing, etc., has rendered this approach impractical. Another method is source code-based log parsing (Nagappan et al., 2009; Xu et al., 2009), which is also challenging as third-party libraries are widely used and some library source codes may not be accessible.

### v2

@@ Abstract.
Log parsing, a vital task for interpreting the vast and complex data produced within software architectures faces significant challenges in the transition from academic benchmarks to the industrial domain. Existing log parsers, while highly effective on standardized public datasets, struggle to maintain performance and efficiency when confronted with the sheer scale and diversity of real-world industrial logs. These challenges are two-fold: 1) massive log templates: The performance and efficiency of most existing parsers will be significantly reduced when logs of growing quantities and different lengths; 2) Complex and changeable semantics: Traditional template-matching algorithms cannot accurately match the log templates of complicated industrial logs because they cannot utilize cross-language logs with similar semantics. To address these issues, we propose ECLIPSE, Enhanced Cross-Lingual Industrial log Parsing with Semantic Entropy-LCS, since cross-language logs can robustly parse industrial logs. On the one hand, it integrates two efficient data-driven template-matching algorithms and Faiss indexing. On the other hand, driven by the powerful semantic understanding ability of the Large Language Model (LLM), the semantics of log keywords were accurately extracted, and the retrieval space was effectively reduced. Notably, we launch a Chinese and English cross-platform industrial log parsing benchmark ECLIPSE-Bench to evaluate the performance of mainstream parsers in industrial scenarios. Our experimental results across public benchmarks and ECLIPSE-Bench underscore the superior performance and robustness of our proposed ECLIPSE. Notably, ECLIPSE both delivers state-of-the-art performance when compared to strong baselines and preserves a significant edge in processing efficiency [^1] We will release our code and dataset..
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ; footnotetext: † Corresponding author.
@@ 1. Introduction
Logs play a crucial role in Algorithmic IT Operations (AIOps) behavior compared to other operational and maintenance data, as they provide significant insights into system behavior (He et al., 2021; Zhu et al., 2019). By analyzing logs, we can complete a variety of downstream tasks, such as anomaly detection (Du et al., 2017; Nandi et al., 2016; Zhang et al., 2024a), fault diagnosis (He et al., 2018; Zhang et al., 2024b) and root cause analysis (Tak et al., 2016; Chuah et al., 2010). In general, most log analysis methods make log parsing the primary step in automated log analysis (He et al., 2016). Illustrated in 1, semi-structured logs are generated from log statement code and need to be parsed to structured logs by algorithms. The most challenging task among them is to extract templates from logs (Zhu et al., 2019), which represent the unchanged parts of the log parsing process. Traditional log parsing was accomplished by manually configuring regular expressions (Lang, 2013). With the exponential growth in the volume and complexity of logs, cloud computing has rendered this approach impractical. Another method is source code-based log parsing (Nagappan et al., 2009; Xu et al., 2009), which are widely used as third-party libraries and some library source codes may not be accessible.